# Hypergraph Characterization of Fusion Rings

### v1

@@ 1. Introduction
Fusion categories are often stratified by their rank, the number of isomorphism classes of simple objects. However, even for fixed rank fusion categories are difficult to fully characterize. At the time of writing a full classification is only known through rank 3, though partial results have been obtained up through rank 5 [18, 20, 6, 7, 14]. Part of the underlying classification problem is a lack of structure and enumeration tools. One partial solution is to characterize the Grothendieck rings of fusion categories rather than the categories themselves. This alleviates certain burdens associated with categorification and associativity, however the problem still remains difficult. In fact, there is currently no known algorithm to enumerate all candidate rings. In this paper we take a first step toward resolving this problem. To accomplish this, we will develop a correspondence between fusion rings and certain digraph/hypergraph pairs. This will enable us to apply tools from graph theory and combinatorics, fields accustomed to handling mathematical constructs with little structure. Our techniques will allow us to completely characterize certain multiplicity-free self-dual fusion rings and are extendable to cases with more complex duality structures and rings with multiplicity. Our main result is:
@@ 2.1. Fusion Rings
A fusion category over $\mathbb{C}$, $\mathcal{C}$, is a $\mathbb{C}$-linear semisimple rigid monoidal category with finitely many simple objects and finite dimensional spaces of morphisms [18]. The fusion category is said to be braided if the monoidal product, $\otimes$, is commutative up to isomorphism with that isomorphism subject to certain consistency conditions [3]. We will denote the set of isomorphism classes of simple objects in $\mathcal{C}$ by $\operatorname{Irr}\left(\mathcal{C}\right)$ and the isomorphism classes of simple objects themselves by $X_{a}$. These simple objects will always be indexed from 00 and ordered such that $X_{0}=\mathbbm{1}$ is the monoidal unit. The dual of an object, $X$, is denoted by $X^{*}$ and the category is said to be self-dual if $X_{a}=X_{a}^{*}$ for all $X_{a}\in\operatorname{Irr}\left(\mathcal{C}\right)$. Duality induces an involution on $\operatorname{Irr}\left(\mathcal{C}\right)$ as well as the index set of $\operatorname{Irr}\left(\mathcal{C}\right)$, specifically $a^{*}$ is defined by the relationship $X_{a^{*}}=\left(X_{a}\right)^{*}$. The number of isomorphism classes of simple objects is referred to as the rank of $\mathcal{C}$. The dimensions of the hom-spaces define the fusion matrices, $N_{a}$, by $\left(N_{a}\right)_{b,c}=N_{a,b}^{c}=\dim_{\mathbb{C}}\operatorname{Hom}_{\mathcal{C}}\left(X_{a}\otimes X_{b},X_{c}\right)$. The fusion matrices are normal non-negative matrices and hence are subject to the Frobenius-Perron Theorem. The Frobenius-Perron eigenvalue of $N_{a}$ is called the Frobenius-Perron dimension or FP-dimension of $X_{a}$ and is denoted by $\operatorname{FPdim}\left(X_{a}\right)$ or more succinctly as $d_{a}$. The FP-dimension of $\mathcal{C}$ is defined by $\operatorname{FPdim}\left(\mathcal{C}\right)=\sum_{a=0}^{r-1}d_{a}^{2}$, where $r$ is the rank of the fusion category $\mathcal{C}$.
@@ 2.2. Graphs and Hypergraphs
While our primary motivation is to study fusion rings, we will do so through the lens specifically graph theory. For this reason we will take a short detour and define the essential terminology for graphs and hypergraphs. Once this is accomplished we will establish a correspondence between fusion rings and graph/hypergraph pairs.
A (simple) graph $G$ is a ordered pair $(V,E)$ where $V$ is a finite set of vertices and $E$ is a subset of the 2-elements subsets of $V$ referred to as the edges. We say two vertices are adjacent, denoted $u\sim v$, if $\set{u,v}\in E$. A graph is said to be trivial if it contains exactly one vertex and empty if the edge set is empty. For a vertex $v$, we define its neighborhood as $N(v)=\set{u\mid\set{u,v}\in E}$. The cardinality of the neighborhood of $v$ is the degree of $v$, $\deg(v)$. We will denote by $\delta$ and $\Delta$, the minimum and maximum degree in a graph.
@@
Similarly to a graph, a directed graph or digraph, is an ordered pair $(V,E)$ where $E$ is a subset of $V\times V$. Much of the terminology above extends naturally to digraphs with additional consideration taken for the direction of the edges. For example, instead of the neighborhood of a vertex $v$ we consider the in- and out-neighborhoods of a vertex, $N^{-}(v)=\set{u\mid(u,v)\in E}$ and $N^{+}(v)=\set{u\mid(v,u)\in E}$, respectively. The in- and out-degrees of a vertex are defined analogously and denoted $\deg^{-}(v)$ and $\deg^{+}(v)$. An edge of the form $(v,v)$ is called a loop and we will say a vertex is trivial if it is an isolated vertex without a self-loop.
@@ 3. Fusion Ring/Digraph Correspondence
More concretely, we note that there are $2\binom{r}{2}+r$ possible edges in a $r$-vertex digraph with self-loops and $\binom{r}{3}$ possible edges in a 3-uniform hypergraph on $r$ vertices. Thus we can build a representation of all potential fusion matrices by specifying the edges of a directed graph, $D$, with self-loops on $\set{0,1,\ldots,r-1}$ and the edges of a 3-uniform hypergraph, $H$, on the same set. Specifically, the digraph and hypergraph encode the fusion rules as follows:
  • for $i\in\set{0,1,\ldots,r-1}$, there is a self-loop at $i$ in $D$ if and only if $N_{ii}^{i}=1$,
  • for $\set{i,j}\subseteq\set{0,1,\ldots,r-1}$ there is a directed edge $(i,j)$ in $D$ if and only if $N_{ii}^{j}=N_{ij}^{i}=N_{ji}^{i}=1$, and
  • for $\set{i,j,k}\subseteq\set{0,1,\ldots,r-1}$ there is a hyperedge $\set{i,j,k}$ in $H$ if and only if $N_{ij}^{k}=N_{ji}^{k}=N_{ik}^{j}=N_{ki}^{j}=N_{jk}^{i}=N_{kj}^{i}=1$.
We note that the entries of $N_{i}$ complete describe the edges incident with the node $i$, in particular, there is a self-loop at $i$ if and only if $\left(N_{i}\right)_{ii}=1$, a directed edge $(i,j)$ if and only if $\left(N_{i}\right)_{jj}=1$, a directed edge $(j,i)$ if and only if $\left(N_{i}\right)_{ij}=\left(N_{i}\right)_{ji}=1$, and a hyperedge $\set{i,j,k}$ if and only if $\left(N_{i}\right)_{jk}=\left(N_{i}\right)_{kj}=1.$ In particular, the edges incident with 00 are completely determined by the convention that $X_{0}=\mathbbm{1}$ and thus $N_{0}=I$. Specifically, the only edges incident to 00 are a self-loop and a directed edge from 00 to $i\in[r-1]=\set{1,\ldots,r-1}.$ Thus for the remainder, we will focus on the correspondence between a self-dual, multiplicity-free rank $r$ fusion ring and a digraph/hypergraph pair on $[r-1]$. In order to illustrate the correspondence between fusion rule, fusion matrix, a graphical presentations of a fusion ring, in Figure 1, we present all three presentations for the fusion ring $\operatorname{Rep}\left(\mathbb{Z}_{3}^{2}\rtimes\mathbb{Z}_{2}\right)$.
For an arbitrary digraph, hypergraph pair $(D,H)$ we will say that $(D,H)$ generates a fusion ring if the associated matrices $\set{N_{i}}$ are the fusion matrices for some fusion ring. We will further say that a digraph $D$ generates a fusion ring if there exists some hypergraph $H$ such that the pair $(D,H)$ generates a fusion ring. Finally, we say that $D$ (resp. $(D,H)$) generates a fusion category if $D$ (resp. $(D,H)$) generates the Grothendieck ring of the category.
@@ 4. Fusion Rings Generated by Undirected Graphs
$w_{ij}$ $\equiv\textrm{number of $k$ such that $\set{i,j,k}$ is a hyperedge.}$.
@@
Combinatorially, this equation can be interpreted as saying that if $i$ and $j$ are adjacent, then the number of hyperedges is one less than the number of triangles containing $\set{i,j}$ in $G$ plus the number of loops on $i$ or $j$. If $i$ and $j$ are not adjacent, then the number of hyperedges containing $\set{i,j}$ is one more than the number of common neighbors of $i$ and $j$ in $G$.
@@ 4.1. Empty Graphs
Let $G$ be the empty graph on $n$ vertices and let $H$ be such that $(G,H)$ generates a fusion ring, then $n=2^{k}-1$ for some $k$ and $H$ is a Steiner triple system of order $12^{k}-$ 1. Furthermore, the $H$ that generates the fusion ring is unique up to isomorphism.
@@
Since $G$ is an empty graph, for any $1\leq i<j\leq n$ we have that $N_{ii}^{j}=N_{jj}^{i}=0$. As a consequence, for all such $i$ and $j$, $(N_{i}N_{j})_{ij}=\sum_{k}N_{ii}^{k}N_{jk}^{j}=N_{ii}^{0}N_{j0}^{0}=1$. Observing that by the commutativity we have that that $1=(N_{j}N_{i})_{ij}=\sum_{k}N_{ji}^{k}N_{ik}^{j}$ and that this counts the number of hyperedges containing $\set{i,j}$ plus the number of directed edges between $i$ and $j$, we conclude that there is a unique hyperedge containing $\set{i,j}$. That is, $H=([n],E)$ is a Steiner triple system. As Steiner triple systems are known to exist if and only if $n\equiv 1,2\pmod{6}$, (see for instance [15]) this significantly restricts the possible self-dual, multiplicity-free fusion rings that can be generated by the empty graph. Indeed, letting $H$ be an arbitrary Steiner triple system generates a set of valid fusion coefficients which respects rigidity and that the duality operation is an anti-involution. However, not all such Steiner triple systems will resulting in a commuting family of fusion matrices.
In order to understand which Steiner triple systems result in a commuting family of fusion matrices, we define for each $i\in[n]$ the graph $G_{i}$ to be the graph whose adjacency matrix is $N_{i}$. We note the fact that $N_{ia}^{b}=N_{ib}^{a}$ implies that this is $G_{i}$ is a graph rather than a digraph. Observing that $(N_{i})_{i0}=1$ and that for all $i,j\in[n]$ there is a unique $k$ such that $(N_{i})_{jk}=1$, we conclude that each $N_{i}$ has precisely one non-zero entry in each row, and hence $G_{i}$ is a perfect matching. Finally, noting that for each (not necessarily distinct) $j,k\in\set{0,\ldots,n}$ there is precisely one $i\in\set{0,\ldots,n}$ such that $(N_{i})_{jk}=1$ and thus $\sum_{i=0}^{r}N_{i}=J$ where $J$ is the $(n+1)\times(n+1)$ matrix of ones. Now, as $N_{0}=I$ we have that $\sum_{i=1}^{r}N_{i}=J-I$ which is the adjacency matrix of the complete graph on $n+1$ vertices. Thus we have that the $G_{i}$ form a partition of the edges of $K_{n+1}$ into $n$ commuting perfect matchings. By [1, Theorem 1], such a decomposition exists if and only if $n+1=2^{k}$ for some $k\in\mathbb{N}$. Further, the pair $(G,H)$ generates a fusion ring if and only if $H$ is Steiner triple system defined from a commuting collection of perfect matchings.
@@
By way of induction suppose that $M_{1},\ldots,M_{t}$ is a collection of commuting perfect matchings on $[2^{k}]$ whose union consists of $2^{k-t}$ disjoint $t$-dimensional hypercubes. Now let $M$ be a perfect matching that commutes with and is disjoint from all of the $M_{i}$. From [1], this implies that $M\cup M_{i}$ is a collection of disjoint 4-cycles for all $i$. Let $S$ be the vertex set of one of the $t$-dimensional hypercubes in $\cup_{i}M_{i}$. We claim that either $M[S]$ has zero edges or $2^{t-1}$ edges. To this end suppose $e$ is an edge in $M[S]$ and without loss of generality label the vertices of $S$ with length $t$ binary strings such that one endpoint of $e$ is the binary 00-string string and for all $i$ the endpoints of the edges in $M_{i}$ differ only in the $i^{\textrm{th}}$ component. Now we note that since $M$ and $M_{i}$ commute, if $e$ is an edge in $M[S]$, then the edge with $i^{\textrm{th}}$ component of both endpoints of $e$ changed is also in $M[S]$. We will abuse notation and denote this edge by $e+e_{i}$. By a similar argument we have that for $j\neq i$, $(e+e_{i})+e_{j}$ is also in $M[S]$. Noting that this operation is commutative, the action of each matching $M_{i}$ is of order two, and that there is a collection of actions which reverse the labels of the endpoints of $e$, we have that $M[S]$ contains at least $2^{t-1}$ edges. In fact, it contains exactly this many as $M$ is a perfect matching and $\left|S\right|=2^{t}$.
Thus, for every hypercube in $\cup M_{i}$, there are at most $12^{t}-t-$ 1 commuting perfect matchings on $[2^{k}]$ containing an edge internal to the hypercube. As there are $2^{k-t}$ such hypercubes, there are at most $2^{k-t}\left(2^{t}-t-1\right)=2^{k}-(t+1)2^{k-t}$ commuting hypercubes which have an edge in any of the hypercubes in $\cup M_{i}$. Since there are at $12^{k}-t-$ 1 unused matchings, there is some matching $M_{t+1}$ which does not contain an edge internal to any hypercube in $\cup M_{i}$. By a similar argument as above, $M_{t+1}\cup\bigcup_{i}M_{i}$ forms $2^{k-t-1}$ hypercubes of dimension $t+1$ on $[2^{k}]$. Repeating this process yields $M_{1},\ldots,M_{k}$, a collection of $k$ commuting perfect matchings on $[2^{k}]$ whose union is a hypercube. The above argument also gives that every matching outside $\set{M_{1},\ldots,M_{t}}$ is completely determined by a single edge, as desired. ∎
@@
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category generated by the empty graph on $n$ vertices, then $n=2^{k}-1$ for some $k\in\mathbb{Z}$ and $\mathcal{C}$ is Grothendieck equivalent to $\operatorname{Rep}\left(\mathbb{Z}_{2}^{k}\right)$.

### v2

@@ Abstract.
Work performed while at Pacific Northwest National Laboratory. Corresponding author.
@@ 1. Introduction
Fusion categories are often stratified by their rank, the number of isomorphism classes of simple objects. However, even for fixed rank fusion categories are difficult to fully characterize. At the time this research was initially presented [7], a full classification was only known through rank 3, though partial results were available up through rank 5 [21, 23, 8, 9, 16]. Recently, there has been significant progress, see for instance the thesis of Vercleyen [26] which provides a complete list of all multiplicity-free fusion categories up to rank 7 and was later extended to rank 9 [25], [2] which classifies all modular data of integral modular fusion categories up to rank 13, and [18] which classifies the Grothendieck rings for complex fusion categories of multiplicity one up to rank 6.
Part of the challenge of the underlying classification problem is a lack of structure and enumeration tools. One partial solution is to characterize the Grothendieck rings of fusion categories rather than the categories themselves. This alleviates certain burdens associated with categorification and associativity, however the problem still remains difficult. In fact, at the time this research [7], there were no known algorithms to enumerate all candidate rings. Recently, Vercleyen and Slingerland [25] have been able to generate an exhaustive list of fusion rings for several multiplicities up to rank 9 by using clever search heuristics to reduce the search space. While their computational results subsume the computational results described in this work, we believe that the novel correspondence between fusion rings and certain digraph/hypergraph pairs provides independent value. Indeed, as we illustrate in Section 4, this correspondence will enable us to apply tools from graph theory and combinatorics, fields accustomed to handling mathematical constructs with little structure, to the enumeration problem significantly reducing the enumeration space. Our techniques will allow us to completely characterize certain multiplicity-free self-dual fusion rings and are extendable to cases with more complex duality structures and rings with multiplicity. Our main result is:
@@ 2.1. Fusion Rings
A fusion category over $\mathbb{C}$, $\mathcal{C}$, is a $\mathbb{C}$-linear semisimple rigid monoidal category with finitely many simple objects and finite dimensional spaces of morphisms [21]. The fusion category is said to be braided if the monoidal product, $\otimes$, is commutative up to isomorphism with that isomorphism subject to certain consistency conditions [4]. We will denote the set of isomorphism classes of simple objects in $\mathcal{C}$ by $\operatorname{Irr}\left(\mathcal{C}\right)$ and the isomorphism classes of simple objects themselves by $X_{a}$. These simple objects will always be indexed from 0 and ordered such that $X_{0}=\mathbbm{1}$ is the monoidal unit. For convenience of notation we will denote by $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$ as the non-identity elements of $\operatorname{Irr}\left(\mathcal{C}\right)$, that is $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)=\operatorname{Irr}\left(C\right)\backslash\mathbbm{1}$. The dual of an object, $X$, is denoted by $X^{*}$ and the category is said to be self-dual if $X_{a}=X_{a}^{*}$ for all $X_{a}\in\operatorname{Irr}\left(\mathcal{C}\right)$. Duality induces an involution on $\operatorname{Irr}\left(\mathcal{C}\right)$ as well as the index set of $\operatorname{Irr}\left(\mathcal{C}\right)$, specifically $a^{*}$ is defined by the relationship $X_{a^{*}}=\left(X_{a}\right)^{*}$. The number of isomorphism classes of simple objects is referred to as the rank of $\mathcal{C}$. The dimensions of the hom-spaces define the fusion matrices, $N_{a}$, by $\left(N_{a}\right)_{b,c}=N_{a,b}^{c}=\dim_{\mathbb{C}}\operatorname{Hom}_{\mathcal{C}}\left(X_{a}\otimes X_{b},X_{c}\right)$. The fusion matrices are normal non-negative matrices and hence are subject to the Frobenius-Perron Theorem. The Frobenius-Perron eigenvalue of $N_{a}$ is called the Frobenius-Perron dimension or FP-dimension of $X_{a}$ and is denoted by $\operatorname{FPdim}\left(X_{a}\right)$ or more succinctly as $d_{a}$. The FP-dimension of $\mathcal{C}$ is defined by $\operatorname{FPdim}\left(\mathcal{C}\right)=\sum_{a=0}^{r-1}d_{a}^{2}$, where $r$ is the rank of the fusion category $\mathcal{C}$.
@@ 2.2. Graphs and Hypergraphs
While our primary motivation is to study fusion rings, we will do so through the lens of graph theory. For this reason we will take a short detour and define the essential terminology for graphs and hypergraphs. Once this is accomplished we will establish a correspondence between self-dual, multiplicity-free fusion rings and graph/hypergraph pairs.
A (simple) graph $G$ is a ordered pair $(V,E)$ where $V$ is a finite set of vertices and $E$ is a subset of the 2-elements subsets of $V$ referred to as the edges. We say two vertices are adjacent, denoted $u\sim v$, if $\{u,v\}\in E$. A graph is said to be trivial if it contains exactly one vertex and empty if the edge set is empty. For a vertex $v$, we define its neighborhood as $N(v)=\{u\mid\{u,v\}\in E\}$. The cardinality of the neighborhood of $v$ is the degree of $v$, $\deg(v)$. We will denote by $\delta$ and $\Delta$, the minimum and maximum degree in a graph.
@@
Similarly to a graph, a directed graph or digraph, is an ordered pair $(V,E)$ where $E$ is a subset of $V\times V$. Much of the terminology above extends naturally to digraphs with additional consideration taken for the direction of the edges. For example, instead of the neighborhood of a vertex $v$ we consider the in- and out-neighborhoods of a vertex, $N^{-}(v)=\{u\mid(u,v)\in E\}$ and $N^{+}(v)=\{u\mid(v,u)\in E\}$, respectively. The in- and out-degrees of a vertex are defined analogously and denoted $\deg^{-}(v)$ and $\deg^{+}(v)$. An edge of the form $(v,v)$ is called a loop and we will say a vertex is trivial if it is an isolated vertex without a self-loop.
@@ 3. Fusion Ring/Digraph Correspondence
More concretely, we note that there are $2\binom{r}{2}+r$ possible edges in a $r$-vertex digraph with self-loops and $\binom{r}{3}$ possible edges in a 3-uniform hypergraph on $r$ vertices. Thus we can build a representation of all potential fusion matrices by specifying the edges of a directed graph, $D$, with self-loops on $\{0,1,\ldots,r-1\}$ and the edges of a 3-uniform hypergraph, $H$, on the same set. Specifically, the digraph and hypergraph encode the fusion rules as follows:
  • for $i\in\{0,1,\ldots,r-1\}$, there is a self-loop at $i$ in $D$ if and only if $N_{ii}^{i}=1$,
  • for $\{i,j\}\subseteq\{0,1,\ldots,r-1\}$ there is a directed edge $(i,j)$ in $D$ if and only if $N_{ii}^{j}=N_{ij}^{i}=N_{ji}^{i}=1$, and
  • for $\{i,j,k\}\subseteq\{0,1,\ldots,r-1\}$ there is a hyperedge $\{i,j,k\}$ in $H$ if and only if $N_{ij}^{k}=N_{ji}^{k}=N_{ik}^{j}=N_{ki}^{j}=N_{jk}^{i}=N_{kj}^{i}=1$.
We note that the entries of $N_{i}$ complete describe the edges incident with the node $i$, in particular, there is a self-loop at $i$ if and only if $\left(N_{i}\right)_{ii}=1$, a directed edge $(i,j)$ if and only if $\left(N_{i}\right)_{jj}=1$, a directed edge $(j,i)$ if and only if $\left(N_{i}\right)_{ij}=\left(N_{i}\right)_{ji}=1$, and a hyperedge $\{i,j,k\}$ if and only if $\left(N_{i}\right)_{jk}=\left(N_{i}\right)_{kj}=1.$ In particular, the edges incident with 0 are completely determined by the convention that $X_{0}=\mathbbm{1}$ and thus $N_{0}=I$. Specifically, the only edges incident to 0 are a self-loop and a directed edge from 0 to $i\in[r-1]=\{1,\ldots,r-1\}.$ Thus for the remainder, we will focus on the correspondence between a self-dual, multiplicity-free rank $r$ fusion ring and a digraph/hypergraph pair on [r1][r-1]. In order to illustrate the correspondence between fusion rule, fusion matrix, a graphical presentations of a fusion ring, in Figure 1, we present all three presentations for the fusion ring $\operatorname{Rep}\left(\mathbb{Z}_{3}^{2}\rtimes\mathbb{Z}_{2}\right)$.
For an arbitrary digraph, hypergraph pair $(D,H)$ we will say that $(D,H)$ generates a fusion ring if the associated matrices $\{N_{i}\}$ are the fusion matrices for some fusion ring. We will further say that a digraph $D$ generates a fusion ring if there exists some hypergraph $H$ such that the pair $(D,H)$ generates a fusion ring. Finally, we say that $D$ (resp. $(D,H)$) generates a fusion category if $D$ (resp. $(D,H)$) generates the Grothendieck ring of the category.
Remark.
It is worth noting that the graphical presentation of a self-dual multiplicity-free fusion ring fully captures the invertibility of the non-identity elements. Specifically, if $X_{i}\in\operatorname{Irr}(\mathcal{C})$ is invertible element in $\mathcal{C}$, then, by the self-duality, it is its own inverse, and in particular, $N_{ii}^{j}=0$ for all $j\neq 0$. As a consequence, if $X_{i}$ is not the identity, then $i$ does not have a self-loop or an out-neighbor in the graphical presentation of $\mathcal{C}$. The converse direction follows along similar lines.
@@ 4. Fusion Rings Generated by Undirected Graphs
[…]
@@
Combinatorially, this equation can be interpreted as saying that if $i$ and $j$ are adjacent, then the number of hyperedges is one less than the number of triangles containing $\{i,j\}$ in $G$ plus the number of loops on $i$ or $j$. If $i$ and $j$ are not adjacent, then the number of hyperedges containing $\{i,j\}$ is one more than the number of common neighbors of $i$ and $j$ in $G$.
@@ 4.1. Empty Graphs
Let $G$ be the empty graph on $n$ vertices and let $H$ be such that $(G,H)$ generates a fusion ring, then $n=2^{k}-1$ for some $k$ and $H$ is a Steiner triple system of order $2^{k}-1$. Furthermore, the $H$ that generates the fusion ring is unique up to isomorphism.
@@
Since $G$ is an empty graph, for any $1\leq i<j\leq n$ we have that $N_{ii}^{j}=N_{jj}^{i}=0$. As a consequence, for all such $i$ and $j$, $(N_{i}N_{j})_{ij}=\sum_{k}N_{ii}^{k}N_{jk}^{j}=N_{ii}^{0}N_{j0}^{0}=1$. Observing that by the commutativity we have that that $1=(N_{j}N_{i})_{ij}=\sum_{k}N_{ji}^{k}N_{ik}^{j}$ and that this counts the number of hyperedges containing $\{i,j\}$ plus the number of directed edges between $i$ and $j$, we conclude that there is a unique hyperedge containing $\{i,j\}$. That is, $H=([n],E)$ is a Steiner triple system. As Steiner triple systems are known to exist if and only if $n\equiv 1,2\pmod{6}$, (see for instance [17]) this significantly restricts the possible self-dual, multiplicity-free fusion rings that can be generated by the empty graph. Indeed, letting $H$ be an arbitrary Steiner triple system generates a set of valid fusion coefficients which respects rigidity and that the duality operation is an anti-involution. However, not all such Steiner triple systems will resulting in a commuting family of fusion matrices.
In order to understand which Steiner triple systems result in a commuting family of fusion matrices, we define for each $i\in[n]$ the graph $G_{i}$ to be the graph whose adjacency matrix is $N_{i}$. We note the fact that $N_{ia}^{b}=N_{ib}^{a}$ implies that this is $G_{i}$ is a graph rather than a digraph. Observing that $(N_{i})_{i0}=1$ and that for all $i,j\in[n]$ there is a unique $k$ such that $(N_{i})_{jk}=1$, we conclude that each $N_{i}$ has precisely one non-zero entry in each row, and hence $G_{i}$ is a perfect matching. Finally, noting that for each (not necessarily distinct) $j,k\in\{0,\ldots,n\}$ there is precisely one $i\in\{0,\ldots,n\}$ such that $(N_{i})_{jk}=1$ and thus $\sum_{i=0}^{r}N_{i}=J$ where $J$ is the $(n+1)\times(n+1)$ matrix of ones. Now, as $N_{0}=I$ we have that $\sum_{i=1}^{r}N_{i}=J-I$ which is the adjacency matrix of the complete graph on $n+1$ vertices. Thus we have that the $G_{i}$ form a partition of the edges of $K_{n+1}$ into $n$ commuting perfect matchings. By [1, Theorem 1], such a decomposition exists if and only if $n+1=2^{k}$ for some $k\in\mathbb{N}$. Further, the pair $(G,H)$ generates a fusion ring if and only if $H$ is Steiner triple system defined from a commuting collection of perfect matchings.
@@
By way of induction suppose that $M_{1},\ldots,M_{t}$ is a collection of commuting perfect matchings on $2^{k}]$ whose union consists of $2^{k-t}$ disjoint $t$-dimensional hypercubes. Now let $M$ be a perfect matching that commutes with and is disjoint from all of the $M_{i}$. From [1], this implies that $M\cup M_{i}$ is a collection of disjoint 4-cycles for all $i$. Let $S$ be the vertex set of one of the $t$-dimensional hypercubes in $\cup_{i}M_{i}$. We claim that either $M[S]$ has zero edges or $2^{t-1}$ edges. To this end suppose $e$ is an edge in $M[S]$ and without loss of generality label the vertices of $S$ with length $t$ binary strings such that one endpoint of $e$ is the binary 0-string string and for all $i$ the endpoints of the edges in $M_{i}$ differ only in the $i^{\textrm{th}}$ component. Now we note that since $M$ and $M_{i}$ commute, if $e$ is an edge in $M[S]$, then the edge with $i^{\textrm{th}}$ component of both endpoints of $e$ changed is also in $M[S]$. We will abuse notation and denote this edge by $e+e_{i}$. By a similar argument we have that for $j\neq i$, $(e+e_{i})+e_{j}$ is also in $M[S]$. Noting that this operation is commutative, the action of each matching $M_{i}$ is of order two, and that there is a collection of actions which reverse the labels of the endpoints of $e$, we have that $M[S]$ contains at least $2^{t-1}$ edges. In fact, it contains exactly this many as $M$ is a perfect matching and $\left|S\right|=2^{t}$.
Thus, for every hypercube in $\cup M_{i}$, there are at most $2^{t}-t-1$ commuting perfect matchings on $2^{k}]$ containing an edge internal to the hypercube. As there are $2^{k-t}$ such hypercubes, there are at most $2^{k-t}\left(2^{t}-t-1\right)=2^{k}-(t+1)2^{k-t}$ commuting hypercubes which have an edge in any of the hypercubes in $\cup M_{i}$. Since there are at $2^{k}-t-1$ unused matchings, there is some matching $M_{t+1}$ which does not contain an edge internal to any hypercube in $\cup M_{i}$. By a similar argument as above, $M_{t+1}\cup\bigcup_{i}M_{i}$ forms $2^{k-t-1}$ hypercubes of dimension $t+1$ on $2^{k}]$. Repeating this process yields $M_{1},\ldots,M_{k}$, a collection of $k$ commuting perfect matchings on $2^{k}]$ whose union is a hypercube. The above argument also gives that every matching outside $\{M_{1},\ldots,M_{t}\}$ is completely determined by a single edge, as desired. ∎
@@
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category of rank $n+1$ such that every element is invertible, then $n=2^{k}-1$ for some $k\in\mathbb{Z}$ and $\mathcal{C}$ is Grothendieck equivalent to $\operatorname{Rep}\left(\mathbb{Z}_{2}^{k}\right)$.